OpenStudy (sbentleyaz):

A particle is moving along the x-axis so that its position at t ≥ 0 is given by s(t) = (t)In(2t). Find the acceleration of the particle when the velocity is first zero.

OpenStudy (solomonzelman):

\(\color{black}{\rm Example}\): A particle is moving along the \(\color{black}{\displaystyle x}\)-axis so that its position (in meters) at \(\color{black}{\displaystyle t\ge0 }\) is given by \(\color{black}{\displaystyle s(t)=t^4-12t }\) (in seconds). Find the acceleration of the particle when the velocity is 20m/s. \(\color{black}{\rm Solution}\): \(\color{black}{\displaystyle s(t)=t^4-12t }\) \(\color{black}{\displaystyle s'(t)=4t^3-12}\) (or \(\color{black}{\displaystyle v(t)=4t^3-12}\)). When (at what time) is the velocity 8m/s? \(\color{black}{\displaystyle v(t)=20}\) \(\color{black}{\displaystyle 4t^3-12=20}\) \(\color{black}{\displaystyle 4t^3=32}\) \(\color{black}{\displaystyle t^3=8}\) \(\color{black}{\displaystyle t=2}\). What is the function for acceleration? \(\color{black}{\displaystyle s'(t)=4t^3-12}\) \(\color{black}{\displaystyle s''(t)=12t^2}\) (or \(\color{black}{\displaystyle a(t)=12t^2}\)). So, at this time (at t=2) what is the acceleration? \(\color{black}{\displaystyle a(2)=12(2)^2=48}\) (m/s\(^2\))

OpenStudy (solomonzelman):

Correction: When (at what time) is the velocity \(\color{red}{20}\)m/s? \(\small \color{black}{(20 ~~{\rm instead~~of~~} 8)}\)